OpenStudy (anonymous):

Write the following equation in slope-intercept form. 3x - 2y = 5

OpenStudy (anonymous):

Well, slope-intercept form is y=mx+b. So we want to isolate the "y" value. To do this, you must move the 2y to the other side, which makes it positive. 3x = 5 +2y Now, move the 5 over to make it negative. 3x - 5 = 2y We want to get 'y' alone, so divide ALL terms by 2: 3/2x - 5/2 = y Therefore, the answer is: \[y = \frac{ 3 }{ 2 }x - \frac{ 5 }{ 2}\]

OpenStudy (anonymous):

Thanks for helping me (:

OpenStudy (anonymous):

No problem!